OpenStudy (anonymous):

i got a 13.5 out of twenty on a test what is that ?

OpenStudy (anonymous):

well, u gotta think, 10 would be 50% so my guess would be about 65 to 75.... im only guessing tho...

OpenStudy (anonymous):

>.<

OpenStudy (loganator):

This has happened to me before you cant always trust answers on this site

OpenStudy (anonymous):

ik it sukx :( but did i at least get a 70

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i need a 70 or higer

OpenStudy (loganator):

You need to check if it makes sense to you as well as other people

OpenStudy (loganator):

thats good :)

OpenStudy (anonymous):

so no one can give me a percentage amount ? lol htats what i really need

OpenStudy (anonymous):

A 70 does sound more reasonable.... to me

OpenStudy (loganator):

Oh how many questions were there and how many did you get wrong

OpenStudy (anonymous):

there were 20

OpenStudy (loganator):

how many wrong

OpenStudy (anonymous):

idk like 6.5 something like that

OpenStudy (anonymous):

how many you answerd after asking here...?/

OpenStudy (anonymous):

all of them since yesterday

OpenStudy (loganator):

yup you got a 70 exactly

OpenStudy (dan815):

13.5/20 = 12/20 + 1.5/20 = 6/10 + 0.75/10 = 67.5%

OpenStudy (anonymous):

NO dan dont say that ^ :'(

OpenStudy (loganator):

this must be a practice test most real tests have 25 questions

OpenStudy (anonymous):

it is a practice test but i need a 70% or higher to take the test

OpenStudy (loganator):

yup, years with flvs, I know lol

OpenStudy (dan815):